### v1

@@ 1 Introduction
The rising frequency and growing severity of natural disasters, such as earthquakes, floods, and hurricanes, cause a significant socio-economic disturbance on a global scale [1]. These events have a devastating impact on the economy, human health, and livelihood. In 2022 alone, there were 387 natural disasters, resulting in 30,704 deaths and affecting approximately 185 million people, causing millions of dollars of damage costs as reported by the Emergency Event Database (EM-DAT) [^1] https://reliefweb.int/report/world/2022-disasters-numbers. In addition to physical damages, these disasters often lead to homelessness, mental health issues, unemployment, emotional distress, financial difficulties, and economic recession. Therefore, it is crucial to consistently explore innovative and enhanced approaches for response and recovery during disasters.
@@
Social media platforms have emerged as invaluable real-time information sources during disaster events [2]. These platforms serve as channels for the rapid dissemination of information, providing critical real-time updates and firsthand accounts during such events [3, 4]. For instance, during the 2023 Turkey-Syria earthquake, social media [^2] https://news.northeastern.edu/2023/02/08/social-media-turkey-earthquake/ served as an indispensable communication platform, facilitating rescue requests and streamlining the coordination of aid efforts. Numerous individuals who were stuck beneath debris utilized mobile tools to access social media platforms, seeking rescue assistance by sharing messages on social media [^3] https://www.theguardian.com/world/video/2023/feb/07/turkey-earthquake-victim-rescued-after-social-media-plea-under-rubble-video, which signifies the utility of social media during such events. However, the research community faces numerous challenges in utilizing social media information during crisis events.
@@
The real-time information provided by social media platforms enhances crisis management capabilities [5], but the overwhelming influx of irrelevant data delays disaster response and management [6] as manual segregation of the relevant information is a time-consuming and costly process. Identifying relevant content amidst this digital noise is a significant challenge. Another important aspect is the incorporation of a diverse mixture of multimodal data [7, 8], encompassing images and text. Each of these modalities presents its own complexities. For instance, the text in crisis-related social media posts often uses informal language, contains frequent misspellings, incorporates numerous hashtags, and mentions other users. Relying solely on text analysis may not capture the full context of a situation [9]. Conversely, images, when viewed in isolation, may lack the necessary context for effective decision-making. To this end, we propose CrisisSpot, a robust and versatile disaster content classification method. The proposed approach merges linguistic context with visual evidence to create a more comprehensive understanding of disaster situations. A system like CrisisSpot empowers stakeholders by allowing them to effectively classify and respond to the most pertinent information contained within social media data posted during a disaster.
@@
It is worth noting that previous works [10, 11, 9, 12, 13] have often overlooked the utilization of Social Context Features (SCF). SCF [14] is an umbrella term that consists of social interaction metrics and semantic text information. They provide a multifaceted view of disaster-related content by incorporating factors such as the credibility of information [15], emotional context [16], and user engagement. In the existing works, various approaches have been proposed over time [17, 9, 12, 18, 19, 13]. Alam et al. [13] utilize the K-nearest neighbor method for constructing a nearest-neighbor graph. In their methodology, nodes are updated based on intra-cluster information, while CrisisSpot updates nodes using neighborhood information selected from the entire dataset across multiple iterations. Consequently, CrisisSpot provides a comprehensive overview of data to each node for feature refinement. Additionally, their choice of neighborhood based on Euclidean distance between nodes is suboptimal for high-dimensional data spaces [20]. To this end, CrisisSpot employs cosine similarity for neighbor selection, a metric proven to perform better in high-dimensional spaces.
Abavisani et al. [18] utilize SSE-Graph [21], which relies on the knowledge graph for swapping embeddings, but it may not explicitly capture the complex structural relationships within the graph. Also, depending on the size of the knowledge graph and the frequency of swapping, SSE-Graph may face scalability challenges. Handling large-scale graphs with frequent embedding swaps might lead to increased computational complexity. To this end, CrisisSpot excels at learning representations that capture the local graph topology, which can be crucial for certain tasks. Additionally, our graph-based approach is designed for scalable graph learning and may provide more efficient solutions for large graphs [22].
The aforementioned studies [9, 12, 18] employ conventional attention mechanisms such as cross-attention to focus solely on aligned information. In contrast, our proposed attention mechanism, Inverted Dual Embedded Attention (IDEA), adeptly directs attention to both aligned and misaligned information patterns across diverse modalities. This explicit recognition significantly augments the model’s capability to effectively handle multimodal data. In conclusion, while existing multimodal paradigms have made significant strides, our study highlights some important research gaps. Our work addresses these limitations by integrating multimodal features, acknowledging the importance of Social Context Features (SCF), and introducing the Inverted Dual Embedded Attention (IDEA) mechanism.
@@
    To the best of our knowledge, we are the first to leverage the potential of Social Context Features (SCF) for disaster content classification in multimodal settings. SCF, effectively merging text-centric and user-centric information, elevates model performance by considering content within a broader context encompassing user history, social interactions, and user credibility.
@@
    We conduct comprehensive experiments on two data-sets, such as CrisisMMD and TSEqD, to evaluate the effectiveness of our proposed approach. The experimental outcomes clearly indicate that our approach surpasses state-of-the-art methods.
The subsequent sections of this article are organized as follows: In section 2, we provide a comprehensive review of the relevant literature. section 3 outlines our methodology, while section 4 presents the outcomes of our experimental assessments. Lastly, in section 5, we provide our concluding remarks on this study.
@@ 2.1.1 Disaster Text Classification
Most of the previous research [23] on classifying disaster-related content has primarily concentrated on the textual modality. For instance, Ghafarian et al. [24] propose a method to identify informative tweets, leveraging the assumption that each tweet represents a word distribution. They performed experiments on 20 crisis incidents of various types. Rudra et al. [25] employed vocabulary-agnostic features, including syntactic features and basic lexical elements, to classify tweets as either conveying situational or non-situational information. They conducted experiments using various disaster datasets and compared their approach with the Bag-Of-Words (BOW) model for Hindi and English tweets. Kumar et al. [26] evaluated multiple traditional Machine Learning and advanced Deep Learning algorithms for categorizing tweets related to disasters across six distinct classes. The effectiveness of the models was assessed through testing on four specific disaster scenarios, such as wildfires, earthquakes, hurricanes, and floods. Xie et al. [27] uses a supervised contrastive learning technique for text classification for humanitarian content in social media. They conducted experiments on three distinct datasets, demonstrating that their method led to enhanced classification accuracy for the model. Nevertheless, prior approaches have made significant strides in addressing challenges associated with disaster text classification. However, they fail to address issues inherent to textual data, including non-standard English, spelling and grammatical errors, and unconventional abbreviations. Additionally, these methods neglect crucial informative factors, such as user credibility, user engagement data, and the emotional context within the text. These factors play a vital role in enhancing the model’s classification capabilities, enabling it to identify relevant, reliable, and information-rich content within the text.
@@ 2.1.2 Disaster Image Classification
Due to the limited number of disaster image datasets, the predominant focus has shifted towards the development of comprehensive and high-quality image datasets [28, 29] within the field of disaster management. Most of the research focuses on categorizing the images based on the severity of the damage. Chaudhuri et al. [30] collect geotagged images from regions impacted by earthquakes. They apply a deep learning method to categorize these images, with a specific focus on identifying survivors under the rubble. They find that the deep learning method exhibits improved accuracy in image classification compared to conventional machine learning techniques. Nguyen et al. [31] employs a deep CNN to classify disaster-related images into various categories, such as severe, mild, and no-damage classes. Their CNN-based approach surpassed Bag-of-Visual-Words (BoVW) methods, achieving significant F1-scores between 0.67 and 0.89. Alam et al. [32] conducted image classification of disaster-related images sourced from social media to assess the extent of the damage. Their work underscored the significance of domain-specific fine-tuning of deep Convolutional Neural Networks. However, prior methods have made substantial progress in tackling the issues related to classifying disaster images. However, they fail to address issues inherent to image data, such as variations in data quality and the context-dependent nature of interpretation. These elements significantly contribute to augmenting the model’s capabilities in classifying images.
@@ 2.2 Multimodal Disaster Content Classification
In recent times, researchers have introduced multimodal systems that harness both text and images to extract relevant information. Agarwal et al. [33] uses the CrisisMMD dataset [34] to build a multimodal end-to-end gated system that identifies the infrastructural damage cues in the data and further qualifies the damage severity of the data based on the ordinal quantification given in the dataset. They surpass state-of-the-art methods, by a significant amount. Abavisani et al. [18] proposed a new cross-attention module during the data fusion process to avoid propagating of false information. Koshy et al. [35] have introduced an advanced deep learning framework that integrates various input modalities, incorporating text and images derived from user-generated content. Their approach outlines a systematic methodology for distinguishing informative tweets from the vast volume of user-generated content on social media platforms. Madichetty et al. [36] have presented a novel technique for identifying informative tweets with multimodal content during disaster situations. Their method involves the utilization of pre-trained RoBERTa for text features and VGG-16 for image features. To combine these two model outputs, they employ a multiplicative fusion technique. However, prior multimodal methods have made significant progress in handling the issues related to disaster content classification. However, most of the prior methods did not incorporate graph-based learning techniques, which helps us enrich the features. Additionally, they overlooked various social interaction information, which enhances the classification abilities of the model.
@@ 3.1.1 Textual Feature Extraction
Each input modality is subjected to an initial encoding process, generating separate unimodal representations labeled as $H_{t}$ for the textual modality and $H_{v}$ for the visual modality. In the context of text modality, given a sequence of text tokens denoted as $S_{t}=\{T_{1},T_{2},T_{3},\ldots,T_{n}\}$, we employ the BERT-based [37] tokenization process to extract token embeddings, resulting in the matrix $H_{t}$ expressed in the following equation.
@@
In Eq. 1, $H_{t}$ represents a matrix containing token embeddings, where $\mathbb{R}^{d\times dt}$ denotes a matrix with dimensions $(d,dt)$. We extract word-level features from the text encoder with a maximum sequence length of 128 and an embedding dimension of 768.

### v2

@@ 1 Introduction
The rising frequency and growing severity of natural disasters, such as earthquakes, floods, and hurricanes, cause a significant socio-economic disturbance on a global scale (Wang and Sun, 2023). These events have a devastating impact on the economy, human health, and livelihood. In 2022 alone, there were 387 natural disasters, resulting in 30,704 deaths and affecting approximately 185 million people, causing millions of dollars of damage costs as reported by the Emergency Event Database (EM-DAT) [^1] https://reliefweb.int/report/world/2022-disasters-numbers. In addition to physical damages, these disasters often lead to homelessness, mental health issues, unemployment, emotional distress, financial difficulties, and economic recession. Therefore, it is crucial to consistently explore innovative and enhanced approaches for response and recovery during disasters.
@@
Social media platforms have emerged as invaluable real-time information sources during disaster events (Basu et al., 2019). These platforms serve as channels for the rapid dissemination of information, providing critical real-time updates and firsthand accounts during such events (Imran et al., 2015; Li et al., 2020). For instance, during the 2023 Turkey-Syria earthquake, social media [^2] https://news.northeastern.edu/2023/02/08/social-media-turkey-earthquake/ served as an indispensable communication platform, facilitating rescue requests and streamlining the coordination of aid efforts. Numerous individuals who were stuck beneath debris utilized mobile tools to access social media platforms, seeking rescue assistance by sharing messages on social media [^3] https://www.theguardian.com/world/video/2023/feb/07/turkey-earthquake-victim-rescued-after-social-media-plea-under-rubble-video, which signifies the utility of social media during such events. However, the research community faces numerous challenges in utilizing social media information during crisis events.
@@
The real-time information provided by social media platforms enhances crisis management capabilities (Biggers et al., 2023), but the overwhelming influx of irrelevant data delays disaster response and management (Shi et al., 2023) as manual segregation of the relevant information is a time-consuming and costly process. Identifying relevant content amidst this digital noise is a significant challenge. Another important aspect is the incorporation of a diverse mixture of multimodal data (Baltrusaitis et al., 2019; Imran et al., 2020), encompassing images and text. Each of these modalities presents its own complexities. For instance, the text in crisis-related social media posts often uses informal language, contains frequent misspellings, incorporates numerous hashtags, and mentions other users. Relying solely on text analysis may not capture the full context of a situation (Ahmad et al., 2022). Conversely, images, when viewed in isolation, may lack the necessary context for effective decision-making. To this end, we propose CrisisSpot, a robust and versatile disaster content classification method. The proposed approach merges linguistic context with visual evidence to create a more comprehensive understanding of disaster situations. A system like CrisisSpot empowers stakeholders by allowing them to effectively classify and respond to the most pertinent information contained within social media data posted during a disaster.
@@
It is worth noting that previous works (Madichetty and Sridevi, 2021; Wang et al., 2023; Ahmad et al., 2022; Wu et al., 2022; Alam et al., 2018a) have often overlooked the utilization of Social Context Features (SCF). SCF (Rehman et al., 2023) is an umbrella term that consists of social interaction metrics and semantic text information. They provide a multifaceted view of disaster-related content by incorporating factors such as the post-user history (Esposito et al., 2022), emotional context (Zhang et al., 2020), and user engagement. In the existing works, various approaches have been proposed over time (Hao and Wang, 2020; Ahmad et al., 2022; Wu et al., 2022; Alam et al., 2018a; Abavisani et al., 2020; Gao et al., 2020). (Alam et al., 2018a) utilize the K-nearest neighbor method for constructing a nearest-neighbor graph. In their methodology, nodes are updated based on intra-cluster information, while CrisisSpot updates nodes using neighborhood information selected from the entire dataset across multiple iterations. Consequently, CrisisSpot provides a comprehensive overview of data to each node for feature refinement. Additionally, their choice of neighborhood based on Euclidean distance between nodes is suboptimal for high-dimensional data spaces (Shirkhorshidi et al., 2015). To this end, CrisisSpot employs cosine similarity for neighbor selection, a metric proven to perform better in high-dimensional spaces.
(Abavisani et al., 2020) utilize SSE-Graph (Wu et al., 2019), which relies on the knowledge graph for swapping embeddings, but it may not explicitly capture the complex structural relationships within the graph. Also, depending on the size of the knowledge graph and the frequency of swapping, SSE-Graph may face scalability challenges. Handling large-scale graphs with frequent embedding swaps might lead to increased computational complexity. To this end, CrisisSpot excels at learning representations that capture the local graph topology, which can be crucial for certain tasks. Additionally, our graph-based approach is designed for scalable graph learning and may provide more efficient solutions for large graphs (Hamilton et al., 2017).
The aforementioned studies (Ahmad et al., 2022; Wu et al., 2022; Abavisani et al., 2020) employ conventional attention mechanisms such as cross-attention to focus solely on aligned information. In contrast, our proposed attention mechanism, Inverted Dual Embedded Attention (IDEA), adeptly directs attention to both aligned and misaligned information patterns across diverse modalities. This explicit recognition significantly augments the model’s capability to effectively handle multimodal data. In conclusion, while existing multimodal paradigms have made significant strides, our study highlights some important research gaps. Our work addresses these limitations by integrating multimodal features, acknowledging the importance of Social Context Features (SCF), and introducing the Inverted Dual Embedded Attention (IDEA) mechanism.
@@
    To the best of our knowledge, we are the first to leverage the potential of Social Context Features (SCF) for disaster content classification in multimodal settings. SCF, effectively merging text-centric and user-centric information, elevates model performance by considering content within a broader context encompassing user history, social interactions, and user post informativeness.
@@
    We conduct comprehensive experiments on two datasets, such as CrisisMMD and TSEqD, to evaluate the effectiveness of our proposed approach. The experimental outcomes clearly indicate that our approach surpasses state-of-the-art methods.
The subsequent sections of this article are organized as follows: In section 2, we provide a comprehensive review of the relevant literature.section 3 outlines our problem definition, section 4 outlines methodology, section 5 presents the outcomes of our experimental assessments. section 6 presents discussions. Lastly, in section 7, we provide our concluding remarks on this study.
@@ 2.1.1 Disaster Text Classification
Most of the previous research (Madichetty and M., 2021) on classifying disaster-related content has primarily concentrated on the textual modality. For instance, (Ghafarian and Yazdi, 2020) propose a method to identify informative tweets, leveraging the assumption that each tweet represents a word distribution. They perform experiments on 20 crisis incidents. (Rudra et al., 2018) employ vocabulary-agnostic features, including syntactic features and basic lexical elements, to classify tweets as either conveying situational or non-situational information. They conduct experiments using various disaster datasets and compare their approach with the Bag-Of-Words (BOW) model for Hindi and English tweets. (Kumar et al., 2019) evaluate multiple traditional Machine Learning and advanced Deep Learning algorithms for categorizing tweets related to disasters across six distinct classes. The effectiveness of the model is assessed through testing on four specific disaster scenarios, such as wildfires, earthquakes, hurricanes, and floods. (Xie et al., 2022) propose a supervised contrastive learning framework (SCMC) for multi-level text classification of disaster information in social media. This framework incorporates data reconstruction and an encoder-decoder architecture to effectively handle disaster-related information. They conduct experiments on three distinct datasets, demonstrating that their method leads to enhanced classification performance. (Wei et al., 2024) propose a novel multi-label text classification approach that addresses the limitations arising from imbalanced label distribution and unclear label hierarchies. They introduce the multi-level constraint augmentation method that leverages historical generation information, sample text content, and topic to generate balanced and relevant samples, specifically targeting the issue of imbalanced labels. They also introduce a label association interaction mechanism that combines text information with label weights, allowing the model to capture co-occurrence patterns between labels and effectively distinguish similar labels. Nevertheless, prior approaches have made significant strides in addressing challenges associated with disaster text classification. However, these methods neglect crucial informative factors, such as user post-history, user engagement metrics, and the emotional context within the text. These factors play a vital role in enhancing the classification capabilities of a method, enabling it to identify relevant, reliable, and information-rich content within the text.
@@ 2.1.2 Disaster Image Classification
Due to the limited number of disaster image datasets, the predominant focus has shifted towards the development of comprehensive and high-quality image datasets (Weber et al., 2023; Alam et al., 2023) within the field of disaster management. Most of the research focuses on categorizing the images based on the severity of the damage. (Chaudhuri and Bose, 2020) collect geotagged images from regions impacted by earthquakes. They apply a deep learning method to categorize these images, with a specific focus on identifying survivors under the rubble. They find that the deep learning method exhibits improved accuracy in image classification compared to conventional machine learning techniques. (Nguyen et al., 2017) employ a deep CNN to classify disaster-related images into various categories, such as severe, mild, and no-damage classes. Their CNN-based approach surpassed Bag-of-Visual-Words (BoVW) methods, achieving significant F1-scores between 0.67 and 0.89. (Alam et al., 2020) conduct image classification of disaster-related images sourced from social media to assess the extent of the damage. Their work underscores the significance of domain-specific fine-tuning of deep Convolutional Neural Networks. The existing methods have made substantial progress in tackling the issues related to classifying disaster images. However, they fail to address issues inherent to image data, such as variations in data quality and the context-dependent nature of interpretation. These elements significantly contribute to augmenting the model’s capabilities in classifying images.
@@ 2.2 Multimodal Disaster Content Classification
In recent times, researchers have introduced multimodal systems that harness both text and images to extract relevant information. (Agarwal et al., 2020) use the CrisisMMD dataset (Alam et al., 2018b) to build a multimodal end-to-end gated system that identifies the infrastructural damage cues in the data and further qualifies the damage severity of the data based on the ordinal quantification given in the dataset. They surpass state-of-the-art methods by a significant amount. (Abavisani et al., 2020) propose a new cross-attention module during the data fusion process to avoid the propagation of false information. (Koshy and Elango, 2023) introduce an advanced Deep Learning framework that integrates various input modalities, incorporating text and images derived from user-generated content. Their approach outlines a systematic methodology for distinguishing informative tweets from the vast volume of user-generated content on social media platforms. (Madichetty et al., 2023) have presented a novel technique for identifying informative tweets with multimodal content during disaster situations. Their method involves the utilization of pre-trained RoBERTa for text features and VGG-16 for image features. To combine these two model outputs, they employ a multiplicative fusion technique. Prior multimodal methods have made significant progress in handling the issues related to disaster content classification. However, most of the prior methods do not incorporate graph-based learning techniques, which helps us enrich the features. Additionally, they overlook various social interaction information, which enhances the classification abilities of the model.
2.3 Attention Mechanisms
Attention mechanisms are vital components of Deep Learning models, particularly for tasks involving sequential data such as text or image features. They enable models to focus on relevant parts of input data. Classical attention mechanisms include: Additive Attention (Bahdanau et al., 2014), Self Attention (Vaswani et al., 2017), and Cross Attention (Hao et al., 2017). Additive attention utilizes a scoring function to assign weights to elements in a sequence, creating a weighted sum that highlights important information. They are widely used in tasks such as machine translation. Self-Attention focuses on analyzing the relationships within a single sequence and employs a scoring function to calculate attention scores. However, instead of comparing each element to a single vector, self-attention compares each element to all other elements in the sequence. This allows the model to identify long-range dependencies within the data, which can be crucial for tasks such as text summarization. Most of the existing works (Rezk et al., 2023; Yoon et al., 2023; Khattar and Quadri, 2022; Abavisani et al., 2020) use Cross-Attention to analyze relationships between two sequences, useful in tasks such as machine translation or visual question answering. It allows models to leverage information from different modalities for better performance. However, these traditional attention mechanisms primarily focus on identifying supportive information within data. In complex domains such as disaster classification, sentiment analysis, and anomaly detection contrary information is also crucial. Current attention mechanisms struggle to distinguish between situations with aligned information and those with misaligned information. To tackle this, we introduce Inverted Dual Embedded Attention (IDEA). IDEA can identify both supportive and non-supportive information patterns within complex data. It achieves this with two modules: a Harmonious Attention Module (HAM) which focuses on elements reinforcing the main concept and a Contrary Attention Module (CAM) which attends to elements deviating from the main concept, capturing contrary information patterns. IDEA combines outputs from both modules to create a rich representation, enabling accurate differentiation between situations with clear information and those with potentially conflicting details, improving disaster classification.
3 Problem Definition
Let $\mathcal{D}=\{\{T_{i}\},\{I_{i}\}\}_{i=1}^{N}$ be a multimodal dataset containing $N$ number of samples with $T_{i}=\{t_{i}\}_{i=1}^{N}$ represents the textual modality for the $i^{\text{th}}$ sample, where $t_{i}$ is the textual content and $I_{i}=\{i_{i}\}_{i=1}^{N}$ represents the visual modality for the $i^{\text{th}}$ sample and where, $i_{i}$ is the visual content. We are given a post containing text $t$ and image $i$, and our objective is to tackle the following tasks:
Task 1: Informative and Non-Informative Content Classification
Task 1 is to classify the given post into informative and non-informative categories. Let $y_{i}\in\{0,1\}$ be the label for the $i^{\text{th}}$ post, where 0 represents non-informative and 1 represents the informative post. $f(t,i)$ denotes a function representing a model that takes text ($t$) and image ($i$) as inputs and produces a probability score indicating their informativeness. This task can be represented as shown in Equation 1:
$\hat{y}_{i}=\arg\max_{c\in C}P(y_{i}=c\mid t,i)$ (1)
where $\hat{y}_{i}$ is the predicted label for the $i^{\text{th}}$ post, $\arg\max_{c\in C}$ denotes finding the argument (class) that maximizes the probability and $C$ denotes the maximum number of classes, and $P(y_{i}=c|t,i)$ represents the conditional probability of the $i^{\text{th}}$ post belonging to class c given the text and image information.
Task 2: Humanitarian Content Classification
Task 2 is to classify the informative post into any seven critical and potentially actionable humanitarian categories. Let $z_{i}\in\{1,2,...,7\}$ be the subcategory label for the informative $i^{\text{th}}$ post, where $z_{i}$ can belong to any one of the categories: infrastructure and utility damage, injured or dead people, rescue volunteering or donation effort, vehicle damage, affected individuals, missing or found people, and other relevant information.
We can define a similar function $g(t,i)$ that takes text and image information of an informative post and outputs a probability distribution over the seven categories as shown in Equation 2:
$P(z_{i}=k|t,i)\approx g_{k}(t,i)$ (2)
where $P(z_{i}=k|t,i)$ denotes the conditional probability of the informative $i^{\text{th}}$ post belonging to category $k$, $g_{k}(t,i)$ represents the predicted probability for category $k$ given text and image information obtained from the function $g$.
The research objectives are delineated as follows:
  • Assessing the impact of Social Context Features on crisis predictions.
  • Investigating the impact of reconstructed features through Graph Neural Networks on the classification outcomes.
  • To investigate whether the proposed attention mechanism, IDEA, captures aligned and misaligned information patterns.
@@ 4.1.1 Textual Feature Extraction
Each input modality is subjected to an initial encoding process, generating separate unimodal representations labeled as $H_{t}$ for the textual modality and $H_{v}$ for the visual modality. In the context of text modality, given a sequence of text tokens denoted as $S_{t}=\{T_{1},T_{2},T_{3},\ldots,T_{n}\}$, we employ the BERT-based (Devlin et al., 2019) tokenization process to extract token embeddings, resulting in the matrix $H_{t}$ expressed in the following equation.
@@
In Eq.˜3, $H_{t}$ represents a matrix containing token embeddings, where $\mathbb{R}^{d\times dt}$ denotes a matrix with dimensions $(d,dt)$. We extract word-level features from the text encoder with a maximum sequence length of 128 and an embedding dimension of 768.